OpenStudy (anonymous):

How do you do this problem -14=-20y-7x and 10y+4=2x

OpenStudy (anonymous):

depends on what you're trying to solve

OpenStudy (mathstudent55):

You have a system of equations. There are several methods of solving a system of equations. One of the methods is called "substitution." To use substitution, solve one equation for one variable. Then plug that in the other equation to solve for one variable. In your case, I'd use substitution, and I'd start by solving the second equation for x. Can you do that and show what you get? Then we'll continue.

OpenStudy (anonymous):

This is a different type do problem like this one -14=-20y-7x and 10y+4=2x

OpenStudy (anonymous):

-14=-20y-7x +7x +7x 7x-14=20y x=? 7x-14=20(0) 7x-14=0 +14 +14 7x=14 /7 /7 x=2 y=? 7x-14=20y 7(0)-14=20y -14=20y /20 /20 -14/20=y -7/10=y y=-7/10

OpenStudy (anonymous):

Oh ok can you hold on for a minut please

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

I'm having a little trouble understanding can you break it up into parts

OpenStudy (mathstudent55):

The solution above is incorrect. Also, the method above is incorrect.

OpenStudy (anonymous):

Just in case the problem is like this -14=-20y-7x 10y+4=2x

OpenStudy (mathstudent55):

Why don't you read what I wrote above, and try to follow it.

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

Ok I go it

OpenStudy (anonymous):

What's next

OpenStudy (mathstudent55):

You solved the second equation for x? If so, what did you get?

OpenStudy (anonymous):

Can you please hold on